An electron with a rest mass of 9.11×10^-31kg moves in a circular orbit of radius 1.4×10^-7m in a uniform magnetic field of 3.5×10^-1 Tesla, perpendicular to the speed of light with which electrons moves. Find the cyclotron frequency of the moving electron


1
Expert's answer
2020-08-10T19:54:08-0400

According to the special theory of relativity, the angular cyclotron frequency is:


"\omega = \dfrac{eB}{\gamma m_e}"


where "e" and "m_e" are charge and mass of electron respectively, "B" is the magnetic field and "\gamma = \dfrac{1}{\sqrt{1-\left ( \frac{v}{c} \right )^2}}"


If electon moves with the speed of light ( "v = c" ), then, according to the previous expression, "\gamma \to \infty".

Then "\omega =0".


Answer. The cyclotron frequency is 0.